OpenStudy (anonymous):

2x=(360/4) can you explain on how to get the answer correctly? tnx :D

OpenStudy (andriod09):

hey there. ;) First thing is first, what can you do you already know about this equation?

OpenStudy (anonymous):

I just want a further explanation on how they get the answer which is 45? :)

OpenStudy (andriod09):

to get an answer, one must start asking him self more questions that are needed. Do you not? if you can answer the first question, then you will get the equation.

OpenStudy (andriod09):

just so you know, \[2x=(360/4)\]\[\frac {2x}{2}=\frac {90}{2}\] \[x=45\]